OpenStudy (anonymous):

solve by substitution. a) x-2y=4 x-3y=6 i got the solution as (8,16) but i checked the answer at the back of the book and that's not it.

OpenStudy (anonymous):

btw it says the answer is (0,-2)

OpenStudy (goformit100):

Recheck your procedure of solving it.

OpenStudy (anonymous):

what is a .jpglarge filetype?

OpenStudy (anonymous):

https://twitter.com/Osgiliaths/status/332694843333693441/photo/1 ok can you see this?

OpenStudy (anonymous):

perfect, thank you

OpenStudy (anonymous):

Have you tried making 2 into an equation so that X is by itself is on one side then the rest of the equation on the other side?

OpenStudy (texaschic101):

x - 2y = 4---> x = 2y + 4 now sub that in to the other equation x - 3y = 6 2y + 4 - 3y = 6 -y = 6 - 4 -y = 2 y = - 2 now sub -2 in for y in either of the original equations x - 2y = 4 x - 2(-2) = 4 x + 4 = 4 x = 4 - 4 x = 0 ANSWER (0,-2)

OpenStudy (anonymous):

Instead of giving an answer, you could point in the right direction, but yes that is the correct way to solve the question. I just hope you actually understood how it's done rather than copy it like a parrot....

OpenStudy (texaschic101):

@skittikinz ...if you have any questions, please ask them.I will be glad to explain anything I did.

OpenStudy (texaschic101):

Did you find your mistake ?

OpenStudy (anonymous):

OH yes i found it! thank you so much

OpenStudy (texaschic101):

you could have also subbed in x = 3y + 6 in for x. Either way , you will get the same answer.

OpenStudy (texaschic101):

Do you have any questions ?

OpenStudy (anonymous):

nope! thanks again